If 50 mL of a 0.4 M NaCl solution is evaporated to dryness, how many grams of salt will remain?
1
Expert's answer
2019-09-16T04:03:03-0400

n(NaCl)=c*V=0.4M*0.05L=0.02 mol

m(NaCl)=n*M=0.02mol*58.44g/mol = 1.168g